OpenStudy (anonymous):

Which of the following is the solution of log (x+3) + 30.001 = –3 ?

OpenStudy (anonymous):

\[\log(x+3)=-3-30.001\]

OpenStudy (anonymous):

then what?

OpenStudy (mathstudent55):

First, subtract 30.001 from both sides. Then use the definition of log as an exponent to solve.

OpenStudy (anonymous):

\[x+3=10^{-33.001}\]

OpenStudy (anonymous):

\[x=10^{-33.001}-3\]

OpenStudy (mathstudent55):